OpenStudy (anonymous):

solve the differential equation: dy/ dx = (6x^2) / (y(1+x^2))

OpenStudy (anonymous):

\[ydy=\frac{ 6x^2 }{ 1+x^2 }dx\]

OpenStudy (anonymous):

\[\int\limits_{}^{}ydy = 6 \int\limits_{}^{}\frac{ (x^2 +1) -1 }{ 1+x^2 }\]

OpenStudy (anonymous):

do you know how to get the answers in ln? the options are 1. y = lnI1+x^3L +c 2. y^2 = -4lnI1+x^3I +c 3. y^2 = 4 ln I 1+x^2I +k where k = 2c

OpenStudy (anonymous):

\[\int\limits_{}^{} y dy = 6 \int\limits_{}^{} dx - \int\limits_{}^{}\frac{ 1 }{ 1+x^2 } dx\]

OpenStudy (anonymous):

\[y^2 = x - tanx + c\]

OpenStudy (anonymous):

actually i dont know how to solve it in ln, sorry !!

OpenStudy (anonymous):

well what if you derive each answer ??